Fermi National Laboratory
 

Operation Summary for September 15, 2010 - September 17, 2010

 

Wednesday Sept. 15

The day shift began with the Tevatron (TeV) in HEP shot setup, with Operations monitoring the Recycler stash (356.2E10), with the Antiproton Source (Pbar) stacking (26.6E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with MiniBooNE off for an access, with the NuMI experiment taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, RF technicians worked on a TeV RF station (TRF3) and controls personnel replaced the VME crate for the TeV sector D1 BPM system. Once these repairs ended, operators completed their search and secure of the D0 enclosure at 9:41 AM. Operations put the TeV through a dry squeeze, after which the RF techs again worked on TRF3; they completed their work at 11:21 AM. Operations put the TeV through another dry squeeze. After the dry squeeze, the D1 BPMs failed again. Controls personnel investigated and at 3:30 PM decided to reinstall the old VME crate. Operators conducted orbit checks. The MI-30 quadrupole power supply VCB opened up on a volts to ground indication; EE Support personnel traced this to a bad current sensor, which was bypassed.

During the evening shift, Operations put the TeV into shot setup at 4:47 PM and established store 8091 at 5:58 PM with an average initial luminosity (AIL) of 308.1E30. The antiprotons, left in the Recycler after partial mining for shot setup, were exceedingly 'hot' and needed to be cooled before transferring fresh antiprotons in from the Accumulator. At 9:25 PM, MI beam began falling out while ramping. Operators traced this to a ringing in the MI-10 quadrupole power supply voltage. Operations called in EE Support personnel to investigate.

Thursday Sept. 16

The midnight shift began with Operations monitoring store 8091 (122.93E30), the stash (313.1E10), and the Pbar stack (26.8E10), with Pbar in standby while EE Support personnel investigated a power supply problem, with NuMI off due to the MI problem, and with MiniBooNE taking beam.

During the midnight shift, EE Support personnel devised a workaround for the MI quadrupole power supply issue; a more permanent fix will be done later this week. Pbar resumed stacking and NuMI resumed taking beam at 12:40 AM. Operations terminated store 8091 at 2:41 AM, put the TeV into shot setup, and established store 8092 at 4:06 AM with an AIL of 299.2E30.

During the day shift, Operations established beam to the Meson MTest experiments T-992 and T-995. MI personnel conducted studies to further troubleshoot the SBD problems. Operations reported some downtime due a Linac RF station (LRF1) that tripped on a plate breaker and then due to a low-level RF system problem after the breaker was reset.

During the evening shift, Operations disabled the MTest beam at 6 PM. Operations terminated store 8092 at 7:20 PM and put the TeV into shot setup, but the setup was delayed for 2 hours while a problem with the SyncLite mirror positioning was investigated. Operations established store 8093 at 10:43 AM with an AIL of 295.28E30.

Friday Sept. 17

The midnight shift began with Operations monitoring store 8093 (215.32E30) and the stash (107.5E10), with Pbar stacking (6.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations disabled NuMI beam at 7 AM, so NuMI personnel could reverse the polarity of their horn.

The Plans for Friday and the Weekend

The plans for today are to end the store around 9:30 AM and then let cryo system technicians work on the TeV sector C2 wet engine. Operations will resume sending beam to NuMI after the next store goes in. NuMI will take beam for about two hours and then shut down for eight days to investigate their target helium leak. The plans for the rest of today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/17/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory